A 1.0 mol sample of an ideal gas is kept at 0.0^0 C during an expansion from 3.0L TO 10.0L.
how much work is done by the gas during the expansion?
 

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
Answer ::

As, the temperature of gas is fixed.
This process is isothermal.

Now,
work done (W) = nRT ln ( V2/V1)
= (1)(8.31)(273)ln ( 3/10)
= - 2.7 × 10^3 J.

Thanks
?? ?
one year ago
As, the temperature of gas is fixed.
This process is isothermal.
 
 
work done (W) = nRT ln ( Vf/Vi)
= (1)(8.31)(273)ln ( 10/3)
=  2.7 × 10^3 J.